# One-step Generative Diffusion for Realistic Extreme Image Rescaling

### v2

@@ Abstract
Image rescaling aims to learn the optimal low-resolution (LR) image that can be accurately reconstructed to its original high-resolution (HR) counterpart, providing an efficient image processing and storage method for ultra-high definition media. However, extreme downscaling factors pose significant challenges to the upscaling process due to its highly ill-posed nature, causing existing image rescaling methods to struggle in generating semantically correct structures and perceptual friendly textures. In this work, we propose a novel framework called One-Step Image Rescaling Diffusion (OSIRDiff) for extreme image rescaling, which performs rescaling operations in the latent space of a pre-trained autoencoder and effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model. Specifically, OSIRDiff adopts a pseudo-invertible module to establish the bidirectional mapping between the latent features of the HR image and the target-sized LR image. Then, the rescaled features are refined by a pre-trained diffusion model to generate more faithful and visually pleasing details. The entire model is end-to-end trained to enable the diffusion priors to guide the rescaling process. Considering the spatially non-uniform reconstruction quality of the rescaled latent features, we propose a novel time-step alignment strategy, which can adaptively determine the generative strength of the diffusion model based on the degree of latent reconstruction errors. Extensive experiments demonstrate the superiority of OSIRDiff over previous methods in both quantitative and qualitative evaluations. The code will be available at: https://github.com/xxx/xxx.
@@ 1 Introduction
With the explosive growth of ultra high-resolution (HR) images, image rescaling has become essential for enhancing data storage and transmission efficiency [44], optimizing resource usage in real-time applications [26], and ensuring proper display on devices with varying resolutions [25]. During image rescaling, HR images are downscaled to low-resolution (LR) images for storage, transmission, or display. When HR images are needed, an upscaling process is employed to restore the downscaled images to their original resolution. Unlike image super-resolution (SR) task, image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
Figure 1: Visual comparison of 16$\times$ image rescaling results.
@@
Recently published image rescaling methods can be classified into two main categories: encoder-decoder architecture [14, 31] and invertible neural networks [41, 18]. The encoder-decoder architecture employs two neural networks to perform downscaling and upscaling. While, invertible neural networks view image rescaling as a reversible process, using the forward and backward passes of an invertible neural network to simulate the downscaling and upscaling. Despite the significant improvements in reconstruction quality achieved by these methods, they are limited to small rescaling factors, such as 2$\times$ or 4$\times$.
@@
However, with the advancement of ultra HR imaging devices, there is an increasing need for extreme rescaling factors (e.g., 16$\times$ or 32$\times$) to achieve more efficient storage and transmission of ultra HR data [53, 8]. Existing image rescaling methods often struggle to generate sufficient detail under extreme rescaling factors. Therefore, some approaches propose using pre-trained GANs as priors to constrain the solution space of the reconstructed results. GRAIN [53] and BDFlow [15] use pre-trained StyleGAN [10, 11] to generate high-quality upscaling face images. VQIR [39] leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [4] to achieve extreme rescaling on natural images. Although these methods have made significant progress in perceptual metrics, we find that their reconstruction results suffer from incorrect semantics and structure. As shown in Fig. 1, although VQIR generates richer details compared to HCFlow, it still fails to reconstruct accurate text and image patterns.
@@
To address the aforementioned issues, we propose a novel framework named One-Step Image Rescaling Diffusion (OSIRDiff) for extreme image rescaling, which performs image rescaling in the latent space of a pre-trained autoencoder and leverages the natural image priors stored in large text-to-image diffusion models to restore HR images. Specifically, first, we use an autoencoder to rescale the latent code of the HR image to the target size, and employ a set of invertible neural networks (INN) to perform the mapping from the feature space to the pixel space to obtain the LR image. Then, the rescaled features are fed into a pre-trained diffusion model to perform a single denoising step, thereby enhancing their perceptual quality. Finally, the entire model is trained end-to-end to allow the diffusion priors to guide the rescaling process. Due to the spatially non-uniform reconstruction quality of the rescaled latent features, different regions of the image require varying levels of generative capability for restoration. Previous studies have shown that the time-step in diffusion models directly controls the generative capacity of the model [35]. Therefore, we propose a time-step alignment mechanism that predicts the appropriate time-step based on the reconstruction quality of the rescaled latent features, thereby achieving a balance between fidelity and realism. As shown in Fig. 1, our model significantly enhances the overall quality of the rescaled images compared to other methods.
@@
  • We propose a One-Step Image Rescaling Diffusion framework that performs well in global semantic and structural reconstruction for the extreme image rescaling task.
@@
Figure 2: Overview of the proposed One-Step Image Rescaling Diffusion (OSIRDiff). First, the HR images $x$ are encoded to obtain the latent features $z$. These latent features are then rescaled to the target size using the decoupled feature rescaling module (DFRM), producing the LR images $y$ and outputting the rescaled latent features $\hat{z}$. Next, the rescaled latent features $\hat{z}$ are passed into a pre-trained diffusion model to perform a single denoising step for perceptual enhancement to obtain $\hat{z}_{0}$. Finally, the perceptually enhanced latent features $\hat{z}_{0}$ are decoded to obtain the rescaled image $\hat{x}$.
@@ 2.1 Image Rescaling
Recently, the emergence of the pre-trained Stable Diffusion (SD) model [27] has showcased its efficacy in both content generation [47] and image restoration [34]. Despite this progress, to date, no effort has been made to investigate it within the framework of extreme image rescaling, presenting an unexplored avenue with potential for advancing the field.
@@ 2.2 Diffusion Model for Image SR
Here, we briefly introduce the application of diffusion models in a related task, i.e., image SR. Diffusion models promote the development of natural image SR in two main ways: training diffusion models from scratch and leveraging pre-trained SD model. The former is represented by SR3 [28], which achieves perceptual quality comparable to GANs. The latter is represented by StableSR [34], which aims to build a model capable of handling image restoration under any complex degradation in real-world scenarios. However, due to the inherent difficulty of integrating diffusion models into existing encoder-decoder architectures, few work has yet attempted to combine them with extreme image rescaling. In this work, we decompose extreme image rescaling into two sub-tasks: feature rescaling in the latent space and one-step perception enhancement with the pre-trained SD U-Net.
2.3 Diffusion Acceleration
To accelerate the inference process of diffusion models, researchers have proposed various optimization strategies. Fast diffusion samplers [30, 12, 21, 51] effectively reduce the sampling steps required by pre-trained diffusion models, lowering them from the initial thousand steps to as few as 20–50 steps. However, further reductions often lead to substantial performance degradation. Distillation techniques [23, 29, 42, 52, 45, 24] have shown significant promise in accelerating inference. For example, Progressive Distillation [29] minimizes sampling steps to under 10 by training a sequence of student models with progressively increasing stride. Instaflow [20] advances this further, achieving one-step generation using reflow and distillation methods. Recently, score-based distillation approaches, including Distribution Matching Distillation [45] and Variational Score Distillation [24], have utilized pre-trained text-to-image diffusion models as distribution matching losses to enable one-step text-to-image generation. Although these methods have made significant strides, high training costs remain a challenge. Inspired by Diffusion2GAN [9], we propose a low-overhead training approach based on Low-Rank Adaptation (LoRA) [6], achieving one-step acceleration for diffusion models and adaptability for image rescaling tasks.
@@ 3 Methodology
3.1 Overview of OSIRDiff
As illustrated in Fig. 2, OSIRDiff consists of four parts: Latent Encoding, Feature Rescaling in the Latent Space, One-Step Perception Enhancement and Latent Decoding. First, the input HR images $x$ are mapped to the latent features $z$ via the pre-trained VAE encoder $\mathcal{E}$. Subsequently, the proposed Decoupled Feature Rescaling Module (DFRM) rescales $z$ to the target size, outputting LR images $y$ and the rescaled latent features $\hat{z}$. Our aim is to make $\hat{z}$ as close as possible to $z$ so that the rescaled results align with the latent space of the pre-trained VAE, while ensuring that $y$ and the original HR images $x$ share the same content. Then, the rescaled features $\hat{z}$ are fed into a denoising U-Net network to perform a single denoising step to get the perceptually enhanced latent features $\hat{z}_{0}$. Finally, perceptually enhanced latent features $\hat{z}_{0}$ are decoded via the pre-trained VAE decoder $\mathcal{D}$ to obtain the rescaled image $\hat{x}$. To dynamically adjust the generative capability based on the reconstruction quality of the rescaled latent features, we design a time-step alignment strategy, which consists of a time-step prediction module (TPM) and a refiner module.
@@ 3.2 Feature Rescaling in the Latent Space
In this part, we perform rescaling of the HR latent features $z$, obtaining LR images $y$ and rescaled features $\hat{z}$. A straightforward idea is to construct a transformation chain as $y=G_{e}\left(z\right)$, $\hat{z}=G_{d}\left(y\right)$. Here, $G_{e}$ is the encoder used to downscale the input features to the target sized LR images, while $G_{d}$ represents the decoder used to reconstruct the original HR features. Guidance loss $\mathcal{L}_{gui.}$ and reconstruction loss $\mathcal{L}_{rec.}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
@@
However, these two losses are contradictory to each other. When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR problem [38], and little information regarding the downscaling process can be embedded into the LR image. In contrast, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [22], which results in the LR image lacking any visual content. Therefore, directly applying the guidance loss $\mathcal{L}_{gui.}$ on $y$ could adversely impact subsequent upscaling.
@@
$z_{\text{lr}}=G_{e}\left(x,z\right)$ (3)
@@
$\hat{z}=G_{d}\left(z_{\text{lr}}\right)$ (5)
Here, $F$ and $F^{-1}$ represent the forward and backward processes of the INN, respectively. Given the latent variable $z\in R^{\frac{H}{8}\times\frac{W}{8}\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{e}$ first transforms it into a compact representation $z_{\text{lr}}\in R^{\frac{H}{N}\times\frac{W}{N}\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor. Subsequently, a set of INNs perform the forward process to transform $z_{\text{lr}}$ from the feature domain to the pixel domain and obtain the LR image $y\in R^{\frac{H}{N}\times\frac{W}{N}\times 3}$ by quantization. Symmetrically, in the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $z_{\text{lr}}$. Finally, the decoder $G_{d}$ is utilized to obtain the rescaled feature $\hat{z}$.
@@
Training of this part involves two losses: reconstruction loss and guidance loss. Considering the information lost in quantization operations, the bidirectional mapping between $z_{\text{lr}}$ and $y$ is not exactly reversible. To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow y\rightarrow\hat{y}_{lr}\rightarrow\hat{z}$, which can be expressed as the following formula:
$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{d}\left(G_{e}\left(x,z\right)\right)-z\parallel_{1}+\\ &\parallel G_{d}\left(F^{-1}\left(F\left(G_{e}\left(x,z\right)\right)\right)\right)-z\parallel_{1}\end{split}$ (6)
@@
$\mathcal{L}_{\mathrm{gui.}}=\parallel F\left(G_{e}\left(x,z\right)\right)-\mathrm{Bicubic}(x)\parallel_{1}\$ (7)
@@
Figure 4: The rescaling operation leads to spatially non-uniform reconstruction quality of the rescaled latent features, which depends on the content of the image and the rescaling factor.
3.3 One-Step Perception Enhancement
The latent variable $\hat{z}$ obtained in Sec. 3.2 retains sufficient semantic information about the HR image, but lacks adequate texture. Therefore, we are proposed to leverage the strong priors stored in the pre-trained SD model to enhance the perceptual quality of the rescaling results in this part. Considering the original diffusion model requires a multi-step iterative sampling algorithm, this significantly increases inference time and makes it difficult to scale to practical applications. Inspired by recent work on accelerating diffusion models [9], we adopt an one-step inference strategy. Specifically, we can directly feed the rescaled latent variable $\hat{z}$ into the denoising U-Net $\epsilon=\epsilon_{\theta}\left(\hat{z},t\right)$ to perform a single denoising step as follows:

### v3

# Timestep-Aware Diffusion Model for Extreme Image Rescaling

@@ Abstract
Image rescaling aims to learn the optimal low-resolution (LR) image that can be accurately reconstructed to its original high-resolution (HR) counterpart, providing an efficient image processing and storage method for ultra-high definition media. However, extreme downscaling factors pose significant challenges to the upscaling process due to its highly ill-posed nature, causing existing image rescaling methods to struggle in generating semantically correct structures and perceptual friendly textures. In this work, we propose a novel framework called Timestep-Aware Diffusion Model (TADM) for extreme image rescaling, which performs rescaling operations in the latent space of a pre-trained autoencoder and effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model. Specifically, TADM adopts a pseudo-invertible module to establish the bidirectional mapping between the latent features of the HR image and the target-sized LR image. Then, the rescaled latent features are enhanced by a pre-trained diffusion model to generate more faithful details. Considering the spatially non-uniform degradation caused by the rescaling operation, we propose a novel time-step alignment strategy, which can adaptively allocate the generative capacity of the diffusion model based on the quality of the reconstructed latent features. Extensive experiments demonstrate the superiority of TADM over previous methods in both quantitative and qualitative evaluations. The code will be available at: https://github.com/xxx/xxx.
@@ 1 Introduction
With the explosive growth of ultra high-resolution (HR) images, image rescaling has become essential for enhancing data storage and transmission efficiency [38], optimizing resource usage in real-time applications [22], and ensuring proper display on devices with varying resolutions [21]. During image rescaling, HR images are downscaled to low-resolution (LR) images for storage, transmission, or display. When HR images are needed, an upscaling process is employed to restore the downscaled images to their original resolution. Unlike image super-resolution (SR), image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
@@
Recently proposed image rescaling methods can be classified into two main categories: encoder-decoder architecture [12, 24] and invertible neural networks [36, 17]. The encoder-decoder architecture employs two neural networks to perform downscaling and upscaling. While, invertible neural networks view image rescaling as a reversible process, using the forward and backward passes of an invertible neural network to simulate the downscaling and upscaling. Despite the significant improvements in reconstruction quality achieved by these methods, they are limited to small rescaling factors, such as 2$\times$ or 4$\times$.
@@
However, with the advancement of ultra HR imaging devices, there is an increasing need for extreme rescaling factors (e.g., 16$\times$ or 32$\times$) to achieve more efficient storage and transmission of ultra HR data [46, 7]. Existing image rescaling methods often struggle to generate sufficient detail under extreme rescaling factors. Therefore, some approaches propose using pre-trained GANs as priors to constrain the solution space of the reconstructed results. GRAIN [46] and BDFlow [14] use pre-trained StyleGAN [9, 10] to generate high-quality upscaled face images. VQIR [33] leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [3] to achieve extreme rescaling on natural images. Although these methods have made significant progress in perceptual metrics, their reconstruction results still suffer from incorrect semantics and structure.
@@
To address the aforementioned issues, we propose a novel framework named Timestep-Aware Diffusion Model (TADM) for extreme image rescaling, which performs image rescaling in the latent space of a pre-trained autoencoder and leverages the natural image priors stored in large text-to-image diffusion models to restore HR images. Specifically, first, the HR images are encoded by the via the VAE encoder to obtain the latent features. Then, we use an autoencoder to rescale the latent code of the HR image to the target size, and employ a set of invertible neural networks (INN) to perform the mapping from the feature space to the pixel space to obtain the LR image. Next, the rescaled features are fed into a pre-trained diffusion model to perform a single denoising step, thereby enhancing their perceptual quality. Finally, the enhanced features are decoded via the VAE decoder to obtain the rescaled image. Additionally, we find that the rescaled latent features exhibit non-uniform degradation, which can be interpreted as the addition of noise at varying levels. As a result, different images require different time-steps to perform the denoising process. Previous studies have shown that the time-step in diffusion models directly controls the generative capacity of the model [28]. Therefore, we propose a time-step alignment mechanism that predicts the appropriate time-step based on the reconstruction quality of the rescaled latent features, thereby adaptively tuning the generative capacity of the diffusion model.
@@
  • We propose a Timestep-Aware Diffusion Model that performs well in global semantic and structural reconstruction for the extreme image rescaling task.
@@
Figure 1: Overview of the proposed Timestep-Aware Diffusion Model (TADM). First, the HR images $x$ are encoded to obtain the latent features $z$. The latent features are then rescaled to the target size using the decoupled feature rescaling module (DFRM), producing the LR images $y$ and outputting the rescaled latent features $\hat{z}$. Next, the rescaled latent features $\hat{z}$ are passed into a pre-trained diffusion model to perform a single denoising step for perceptual enhancement to obtain $\hat{z}_{0}$. In this process, a time prediction module (TPM) is used to estimate the time step $t$ based on $\hat{z}$ and the predicted $t$ is then fed into both the U-Net and the time scheduler. Finally, the perceptually enhanced latent features $\hat{z}_{0}$ are decoded to obtain the rescaled image $\hat{x}$.
@@ 2.2 Diffusion Model for Image SR
Here, we briefly introduce the application of diffusion models in a related task, i.e., image SR. Diffusion models promote the development of natural image SR in two main ways: training diffusion models from scratch [13, 4, 40] and leveraging pre-trained SD model [19, 39, 35]. The former is represented by SR3 [23], which achieves perceptual quality comparable to GANs. The latter is represented by StableSR [27], which aims to build a model capable of handling image restoration under any complex degradation in real-world scenarios.
However, the above methods typically require dozens or even hundreds of iterations during inference, which significantly increases computational costs. Researchers have begun exploring how to achieve image SR using one-step diffusion models. SinSR [30], based on ResShift [40], derives a deterministic one-step sampling process and employs a consistency preserving loss to distill the prior knowledge from a multi-step teacher network into a one-step student network. A representative work, OSEDiff [34], directly feeds the LQ image into the SD model and employs variational score distillation for regularization. S3Diff [42] proposes a degradation-aware LoRA and utilizes a pre-trained DINO [43] model as a discriminator for adversarial training. InvSR [41] designs a partial noise prediction strategy to provide a dynamic starting sampling point and supports inference with an arbitrary number of sampling steps ranging from one to five.
@@ 3 Methodology
3.1 Overview of TADM
As illustrated in Fig. 1, TADM consists of four parts: Latent Encoding, Feature Rescaling in the Latent Space, Denoising Guided Perception Enhancement and Latent Decoding. First, the input HR images $x$ are mapped to the latent features $z$ via the pre-trained VAE encoder $\mathcal{E}$. Subsequently, the proposed Decoupled Feature Rescaling Module (DFRM) rescales $z$ to the target size, outputting LR images $y$ and the rescaled latent features $\hat{z}$. Then, the rescaled features $\hat{z}$ are fed into a denoising U-Net network to perform a single denoising step to get the perceptually enhanced latent features $\hat{z}_{0}$. Finally, $\hat{z}_{0}$ is decoded via the pre-trained VAE decoder $\mathcal{D}$ to obtain the rescaled image $\hat{x}$. Overall, the latent rescaling process can be regarded as a noise-adding process, while the latent enhancement process can be viewed as a denoising process. To dynamically estimate the ‘noise’ density introduced by latent rescaling, we design a time alignment strategy, which consists of a time-step prediction module (TPM) and a time scheduler module. It can better address the non-uniform degradation caused by rescaling, thereby enabling the dynamic allocation of generative capabilities of pre-trained SD.
@@ 3.2 Feature Rescaling in the Latent Space
In this part, we perform rescaling of the HR latent features $z$, obtaining LR images $y$ and rescaled features $\hat{z}$. A straightforward idea is to construct a transformation chain as $y=G_{\mathrm{e}}\left(z\right)$, $\hat{z}=G_{\mathrm{d}}\left(y\right)$. Here, $G_{\mathrm{e}}$ is the encoder used to downscale the input features to the target sized LR images, while $G_{\mathrm{d}}$ represents the decoder used to reconstruct the original HR features. Guidance loss $\mathcal{L}_{\mathrm{gui.}}$ and reconstruction loss $\mathcal{L}_{\mathrm{rec.}}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
@@
However, these two losses are contradictory to each other. When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR problem [32], embedding little information about the downscaling process into the LR image. In contrast, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [20], causing the LR image to lack meaningful visual content. Therefore, directly applying the guidance loss $\mathcal{L}_{\mathrm{gui.}}$ on $y$ could adversely impact subsequent upscaling.
@@
$z_{\text{lr}}=G_{\mathrm{e}}\left(x,z\right)$ (3)
@@
$\hat{z}=G_{\mathrm{d}}\left(z_{\text{lr}}\right)$ (5)
Here, $F$ and $F^{-1}$ represent the forward and backward processes of the INN, respectively. Given the latent variable $z\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{\mathrm{e}}$ first transforms it into a compact representation $z_{\text{lr}}\in\mathbb{R}^{\frac{H}{N}\times\frac{W}{N}\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor. Subsequently, a set of INNs perform the forward process to transform $z_{\text{lr}}$ from the feature domain to the pixel domain and obtain the LR image $y\in\mathbb{R}^{\frac{H}{N}\times\frac{W}{N}\times 3}$ by quantization. In the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $z_{\text{lr}}$. Finally, the decoder $G_{\mathrm{d}}$ is utilized to obtain the rescaled feature $\hat{z}$.
@@
Training of this part involves two losses: reconstruction loss and guidance loss. Considering the information lost in quantization operations, the bidirectional mapping between $z_{\text{lr}}$ and $y$ is not exactly reversible. To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow y\rightarrow\hat{z}_{\text{lr}}\rightarrow\hat{z}$, which can be expressed as the following formula:
$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{\mathrm{d}}\left(G_{\mathrm{e}}\left(x,z\right)\right)-z\parallel_{1}+\\ &\parallel G_{\mathrm{d}}\left(F^{-1}\left(F\left(G_{\mathrm{e}}\left(x,z\right)\right)\right)\right)-z\parallel_{1}\end{split}$ (6)
@@
$\mathcal{L}_{\mathrm{gui.}}=\parallel F\left(G_{\mathrm{e}}\left(x,z\right)\right)-\mathrm{Bicubic}(x)\parallel_{1}\$ (7)
@@
3.3 Denoising Guided Perception Enhancement
The latent variable $\hat{z}$ obtained in Sec. 3.2 retains sufficient semantic information about the HR image, but lacks adequate texture. Therefore, we propose to leverage the strong priors stored in the pre-trained SD model to enhance the perceptual quality of the rescaling results in this part. Original diffusion model requires a multi-step sampling algorithm, significantly increases inference time and makes it difficult to scale to practical applications. Inspired by recent work on accelerating diffusion models [8], we adopt an one-step inference strategy. Specifically, we can directly feed the rescaled latent variable $\hat{z}$ into the denoising U-Net $\epsilon=\epsilon_{\theta}\left(\hat{z},t\right)$ and perform a single denoising step as follows: